OpenStudy (goldeverything):

Elise has a budget of sixty dollars to spend on back-to-school clothes. She finds some shirts she likes that cost fifteen dollars each. How many shirts, x, can Elise buy? x is less than four x is greater than four x is less than or equal to four x is greater than or equal to four

OpenStudy (johnweldon1993):

A budget of $60 ...so she cannot go over that but she can meet it so we have a less than or equal to sign If each shirt cost $15 each we have \[15x \le 60\] To solve for x....divide both sides of that equation by 15 and you will get \[x \le ?\]

OpenStudy (mathstudent55):

Each shirt costs $15. 1 shirt costs \( $15 \times 1 = 15\) 2 shirts cost \( $15 \times 2 = 30\) 3 shirts cost \( $15 \times 3 = 45\) x shirts cost \( $15 \times x = 15x\) Since she has a total of $60, her total expense must be less than or equal to $60. \(15x \le 60 \) Divide both sides by 15 to find x.